OpenStudy (anonymous):

WILL GIVE MEDALS 12. which is equivalent to 2x/x-7- x+7/x? A. (x-7)/(x(x-7) B.(x+7)/(x(x-7) C.(x^2+49)/(x(x-7) D.(x^2-49)/(x(x-7)

OpenStudy (anonymous):

@campbell_st can you help

OpenStudy (campbell_st):

well you need a common denominator which will be x(x - 7) so the 1st it becomes \[\frac{2x \times x}{x(x -7)} - \frac{x \times (x - 7)}{x(x - 7)}\] you'll need to simplify the numerator... my suggestion is distribute collect like terms then factor...

OpenStudy (anonymous):

so do i combine all the x's

OpenStudy (campbell_st):

like terms... means the 2 things that are x^2... you can subtract them...

OpenStudy (anonymous):

so i get \[\frac{ (x-7) }{ (x(x-7) }\]

OpenStudy (campbell_st):

oops... slight mistake \[\frac{2x \times x}{x(x -7)} - \frac{(x +7)(x -7)}{x(x - 7)}\] which becomes \[\frac{2x^2}{x(x -7)} - \frac{x^2 - 49}{x(x - 7)}\] which simplifies to \[\frac{x^2 + 49}{x(x -7)}\]

OpenStudy (anonymous):

oh ok got it